# Addressing the Non-perturbative Regime of the Quantum Anharmonic Oscillator by Physics-Informed Neural Networks

### v1

@@ VIII Conclusions
In this work, we utilized a novel unsupervised deep-learning-based model to find the eigenvalues and eigenstates for a potential with a plethora of application in physics. To benchmark the performances of our method, we first started by finding the first six stationary states of the harmonic oscillator. The obtained results then served as initialization to solve the anharmonic oscillator and its pure quartic version for growing values of the parameter $\lambda$. Our method is reliable on the full range of $\lambda$ and it allowed us to establish two exponential scalings for the energy levels. The result seems to suggest there is a critical value $\lambda_{c}$ that decreases with the energy for which the perturbative expansion is not longer reliable. This is consistent with the previous results, in fact, the spectrum does not have a constant spacing due to its nonlinear dependence on the occupation number [14]. Therefore, this opens an interesting avenue of research to study if such conclusions can be drawn also for field theories in higher dimensions and in particular if this could be related to the phase transition that could be studied via mapping of the scalar $\lambda\phi^{4}$ in $(3+1)-$dimensions to an Ising model in $d\geq 4$ dimensions. For the machine learning side, since the changes in the wave function as $\lambda$ increases seem to be smooth it is, in principle, possible to use these results to train a neural network that can take both $\lambda$ and $x$ as the input, and outputs the wave function at $x$ and energy for a given eigenstate of the potential with the given $\lambda$.
@@ Appendix A Coefficients of the linear fits
In this Section we report the numerical coefficients of the fits in Figure 7. In Table 6 we report the interpolation coefficients used to draw the blue dashed lines in Figure 7, equivalently in Tables 6 and 6 we report the interpolation coefficients used to draw the red and green dashed lines in Figure 7. That is, the coefficients of the line:

### v2

@@ VIII Conclusions
In this work, we utilized a novel unsupervised deep-learning-based model to find the eigenvalues and eigenstates for a potential with a plethora of applications in physics. To benchmark the performances of our method, we first started by finding the first six stationary states of the harmonic oscillator. The obtained results then served as initialization to solve the anharmonic oscillator and its pure quartic version for growing values of the parameter $\lambda$, where it has been validated with results obtained via the Lagrangian-mesh method. Our method is reliable on the full range of $\lambda$ and it allowed us to establish two exponential scalings for the energy levels. The result seems to suggest there is a critical value $\lambda_{c}$ that decreases with the energy for which the perturbative expansion is no longer reliable. This is consistent with the previous results, in fact, the spectrum does not have a constant spacing due to its nonlinear dependence on the occupation number [14]. Therefore, this opens an interesting avenue of research to study if such conclusions can be drawn also for field theories in higher dimensions and in particular if this could be related to the phase transition that could be studied via mapping of the scalar $\lambda\phi^{4}$ in $(3+1)-$dimensions to an Ising model in $d\geq 4$ dimensions. For the machine learning side, since the changes in the wave function as $\lambda$ increases seem to be smooth it is, in principle, possible to use these results to train a neural network that can take both $\lambda$ and $x$ as the input, and outputs the wave function at $x$ and energy for a given eigenstate of the potential with the given $\lambda$. As a final remark, we like to stress that already established methods could be more efficient in terms of resources, our work paves the way to the applications of PINNs to complex scenario of interest in quantum science and technology, where classical numerical techniques are struggling.
@@ Appendix A Coefficients of the linear fits
In this Section, we report the numerical coefficients of the fits in Figure 10. In Table 8 we report the interpolation coefficients used to draw the blue dashed lines in Figure 10, equivalently in Tables 8 and 8 we report the interpolation coefficients used to draw the red and green dashed lines in Figure 10. That is, the coefficients of the line: